### v1

# Proximality and selflessness for group C*-algebras

Abstract.
We prove that the reduced group $\mathrm{C}^{*}$-algebra of an infinite discrete group having a topologically-free extreme boundary is selfless in the sense of L. Robert.
Key words and phrases:
Extremely proximal actions, extreme boundaries, selfless $\mathrm{C}^{*}$-algebras
1991 Mathematics Subject Classification:
Primary 37B05; Secondary 46L35
The author was partially supported by JSPS KAKENHI Grant Numbers 24K00527, 25H00588, 25H00593
1. Introduction
A remarkable property called selflessness is recently introduced for a $\mathrm{C}^{*}$-algebra by L. Robert ([Ro]) and it quickly attracted a number of researchers’ attention (see, e.g., [AGKEP], [HKER], [RTV], [Vi]) as it implies many important regularity properties such as simplicity, (in the tracial setting) stable rank one, and strict comparison. Let’s say a group $\Gamma$ is $\mathrm{C}^{*}$-selfless if its reduced group $\mathrm{C}^{*}$-algebra is selfless. In this paper, we prove that every infinite countable discrete group with a minimal topologically-free extremely-proximal action is $\mathrm{C}^{*}$-selfless. Note that every $\mathrm{C}^{*}$-selfless group is $\mathrm{C}^{*}$-simple ([Ro]) and whether the converse is also true is an open problem (Problem XC in [STW]). Our result should be compared with Kalantar and Kennedy’s theorem ([KK]) that a discrete group $\Gamma$ is $\mathrm{C}^{*}$-simple if and only if it has a minimal topological-free strongly-proximal action. As the names suggest, strong proximality is weaker than extreme proximality ([Gl]).
Definition.
Let $\Gamma$ be a countable discrete group and $\Gamma\curvearrowright X$ be an action on a compact topological space $X$. The action $\Gamma\curvearrowright X$ (or the $\Gamma$-space $X$) is called an extreme boundary if it is minimal and extremely proximal ([Gl]) in the sense that for every non-empty open subsets $U$ and $V$ of $X$, there is $g\in\Gamma$ such that $g(X\setminus U)\subset V$.
For $x\in X$, we denote by $\Gamma_{x}\coloneq\{g\in\Gamma:gx=x\}$ the stabilizer subgroup at $x$. The action $\Gamma\curvearrowright X$ is said to be topologically-free if the set of points with trivial stabilizer group is dense in $X$. Note that if an extreme boundary $\Gamma\curvearrowright X$ is not topologically-free, then every proper closed subset of $X$ is pointwise fixed by some $g\in\Gamma\setminus\{1\}$.
We say a sequence $(z_{n})_{n}\in\Gamma$ is axial if there is a topologically-free extreme boundary $\Gamma\curvearrowright X$ with distinct points $z_{\pm}\in X$ that satisfies the following two conditions. (1): For every neighborhoods $U_{\pm}$ of $z_{\pm}$ one has $z_{n}(X\setminus U_{-})\subset U_{+}$ eventually, or equivalently that $z_{n}^{-1}(X\setminus U_{+})\subset U_{-}$ eventually. (2): The $\Gamma$-action on $\{z_{\pm}\}$ is free in the sense that $g\{z_{\pm}\}\cap\{z_{\pm}\}\neq\emptyset$ implies $g=1$. Note that every topologically-free extreme boundary $X$ with $|X|>2$ accommodates an axial sequence.
Theorem 1.
An infinite countable discrete group $\Gamma$ having a topologically-free extreme boundary is $\mathrm{C}^{*}$-selfless. More precisely, for any axial sequence $(z_{n})_{n}$ in $\Gamma$ and any free ultrafilter $\mathcal{U}$, the homomorphism
$\Gamma*\mathopen{\langle}z\mathclose{\rangle}\to\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}},$
given by the constant embedding of $\Gamma$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ and $z\mapsto[\lambda(z_{n})]_{n}\in\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$, induces a faithful embedding of the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(G*\mathopen{\langle}z\mathclose{\rangle})$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$.
See, e.g., [Bo], [BIO], [FLMMS], [IO], [JR], [LS], [LBMB] for examples of extreme boundaries. Every group satisfying the conclusion of Theorem 1 is mixed-identity-free ([HO]) and converse holds true if $\Gamma$ has a faithful minimal action of general type on a tree ([FLMMS]), since such an action gives rise to an extreme boundary. In particular, non-degenerate free product groups are $\mathrm{C}^{*}$-selfless. For the case of amalgamated free products and HNN extensions, see [FLMMS] and [IO] for the characterization of topological freeness of the corresponding Bass–Serre tree boundaries. Non-elementarily relatively hyperbolic groups ([Bo]) with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless, as their Bowditch compactifications are topologically-free extreme boundaries. It is not clear whether this generalizes to acylindrically hyperbolic groups, because their hyperbolic boundaries need not be compact.
Notations
A sequence is written as $(x_{n})_{n}$, while the equivalence class it represents in the ultrapower is written by $x_{n}]_{n}$. We use the symbol $z^{\pm}$ for a slightly abusive notation for “$z$ and/or $z^{-1}$” and similarly for $z_{\pm}$. For $\epsilon\in\{\pm\}$, we denote by $\bar{\epsilon}$ the opposite of $\epsilon$.
Acknowledgments
This research was carried out while the author was visiting the Mathematisches Forschungsinstitut Oberwolfach for the workshop “C*-Algebras”, August 4–8, 2025. He acknowledges the kind hospitality and the exciting environment provided by the institute. He thanks G. Patchell for the stimulating talk there and a useful comment on this paper.
2. Trees and covariant representations
We recall the notion of trees and their compactification. A tree is a connected graph without nontrivial cycles. See Section 5.2 and Appendix E in [BO] for more on this theme. Here we consider a countable tree $\mathrm{T}$ of infinite degree. The tree $\mathrm{T}$ is identified with its vertex set. A geodesic ray is an infinite sequence $\omega(0),\omega(1),\ldots$ in $\mathrm{T}$ such that $\omega(n)$ adjacent to $\omega(n-1)$ and $\omega(n)\neq\omega(n-2)$. Geodesic rays $\omega$ and $\omega^{\prime}$ are equivalent if there are $N$ and $N^{\prime}$ such that $\omega(N+n)=\omega^{\prime}(N^{\prime}+n)$ for all $n\geq 0$. The boundary $\partial\mathrm{T}$ is the equivalence classes of geodesic rays and $\bar{\mathrm{T}}\coloneq\mathrm{T}\cup\partial\mathrm{T}$ is the compactification of $\mathrm{T}$ (whose topology will be introduced shortly). We define $f\colon\mathrm{T}\times\bar{\mathrm{T}}\to\mathrm{T}$ as follows. For $s\in\mathrm{T}$, we set $f(s,s)=s$. For distinct $s,t\in\mathrm{T}$, we define $f(t,s)\in\mathrm{T}$ to be the unique point between $t$ and $s$ that is adjacent to $t$. For $t\in\mathrm{T}$ and $\omega\in\partial\mathrm{T}$, we define $f(t,\omega)\coloneq f(t,\omega(n))$ for any $n$ large enough. For each $t\in\mathrm{T}$, we consider a copy $\mathrm{T}_{t}$ of $\mathrm{T}$ and equip $\mathrm{T}_{t}$ a compact topology by taking the one-point compactification $\mathrm{T}\cup\{\infty\}$ of $\mathrm{T}$ and identify the new point $\infty$ with $t$. That is, every sequence in $\mathrm{T}$ that diverges to infinity converges to $t$ in $\mathrm{T}_{t}$. We consider the embedding
$\bar{\mathrm{T}}\ni\omega\mapsto(f(t,\omega))_{t}\in\prod_{t\in\mathrm{T}}\mathrm{T}_{t}$
and equip $\bar{\mathrm{T}}$ with the induced topology. This makes $\bar{\mathrm{T}}$ a second-countable compact topological space (Proposition 5.2.5 in [BO]). Note that if $s\in\mathrm{T}$ and $t_{n}\in\mathrm{T}$ are such that $f(s,t_{n})\to\infty$ in $\mathrm{T}$, then $t_{n}\to s$ in $\bar{\mathrm{T}}$.
Theorem 2.
Let $\Lambda$ be a countable discrete group acting on a countable tree $\mathrm{T}$ and on a unital $\mathrm{C}^{*}$-algebra $A$. Let $\pi\colon\Lambda\ltimes_{\max}A\to\mathbb{B}(\mathcal{H})$ be a covariant representation. Assume that there is a $\Lambda$-equivariant unital positive map from $C(\bar{\mathrm{T}})$ into the center of $A$ and that for every $t]\in\mathrm{T}/\Lambda$, the restriction of $\pi$ to the stabilizer subgroup $\Lambda_{t}\coloneq\{g\in\Lambda:gt=t\}$ is continuous on the reduced crossed product $\Lambda_{t}\ltimes_{\min}A$. Then $\pi$ is continuous on the reduced crossed product $\Lambda\ltimes_{\min}A$
Proof.
Let $\pi\colon\Lambda\ltimes_{\max}A\to\mathbb{B}(\mathcal{H})$ be given. We denote by $\lambda$ the unitary representation of $\Lambda$ on $\ell_{2}\mathrm{T}$ induced by $\Lambda\curvearrowright\mathrm{T}$ and denote by $\sigma$ the $\Lambda$ action on $A$. We claim that $\pi$ is weakly contained in the representation $\tilde{\pi}$ on $\ell_{2}\mathrm{T}\otimes\mathcal{H}$ given by $g\mapsto(\lambda\otimes\pi)(g)$ for $g\in\Lambda$ and $a\mapsto 1\otimes\pi(a)$ for $a\in A$. Indeed, there is a sequence $\zeta_{n}\colon\bar{\mathrm{T}}\to\mathrm{Prob}(\mathrm{T})$ of Borel maps such that $\sup_{x\in\bar{\mathrm{T}}}\|g\zeta_{n}^{x}-\zeta_{n}^{gx}\|_{1}\to 0$ for every $g\in\Lambda$ (Lemma 5.2.6 in [BO]). It follows that $h_{n}^{t}(x)\coloneq\zeta_{n}^{x}(t)$ defines sequences of positive Borel functions on $\bar{\mathrm{T}}$ such that $\sum_{t\in\mathrm{T}}h_{n}^{t}=1$ and
$\|\sum_{t\in\mathrm{T}}|h_{n}^{gt}-\sigma_{g}(h_{n}^{t})|\|_{\infty}\to 0.$
NB: the above sum is an infinite sum of positive Borel functions. Let $\theta$ denote the $\Lambda$-equivariant unital completely positive map from $C(\bar{\mathrm{T}})$ into the center of $A$. The map $\pi\circ\theta$ canonically extends to a normal unital positive map $\Theta$ from the $\mathrm{C}^{*}$-algebra $B(\bar{\mathrm{T}})$ of the bounded Borel functions into $(\pi\circ\theta)(C(\bar{\mathrm{T}}))^{\prime\prime}$. By “normal”, we mean that it sends a bounded pointwise convergent sequence to a strong operator topology convergent sequence. We note that $\Theta$ satisfies $|\Theta(f)|\leq\Theta(|f|)$ because the range of $\Theta$ is commutative. We define isometries $V_{n}\colon\mathcal{H}\to\ell_{2}\mathrm{T}\otimes\mathcal{H}$ by $V_{n}\xi\coloneq\sum_{t\in\mathrm{T}}\delta_{t}\otimes\Theta(h_{n}^{t})^{1/2}\xi$, or $V_{n}$ is the column operator $(\Theta(h_{n}^{t})^{1/2})_{t\in\mathrm{T}}$. Then
$\|V_{n}\pi(g)-(\lambda\otimes\pi)(g)V_{n}\|^{2}=\|\sum_{t}|\Theta(h_{n}^{gt})^{1/2}-\Theta(\sigma_{g}(h_{n}^{t}))^{1/2}|^{2}\|_{\infty}\to 0$
for every $g\in\Lambda$ and $V_{n}\pi(a)=(1\otimes\pi(a))V_{n}$ for $a\in A$. This proves the claim that $\pi$ is weakly contained in $\tilde{\pi}$. Thus continuity of $\pi$ follows from that of $\tilde{\pi}$.
For the continuity of $\tilde{\pi}$, it suffices to consider the continuity of the vector states $\varphi$ associated with the vectors of the form $\delta_{t}\otimes\xi$, since they form a cyclic family. But $\varphi=\psi\circ E$, where $E$ is the conditional expectation from $\Lambda\ltimes_{\max}A$ onto $\Lambda_{t}\ltimes_{\max}A$ and $\psi$ is the vector state associated with $\xi$. (We do not need the fact that the natural embedding $\Lambda_{t}\ltimes_{\max}A\subset\Lambda\ltimes_{\max}A$ is faithful.) Hence the continuity on the reduced crossed product $\Lambda\ltimes_{\min}A$ follows from that for $\Lambda_{t}\ltimes_{\min}A$ for $t\in\mathrm{T}$ one from each $\Lambda$-orbit. ∎
3. A space with “tree structure”
Let $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$. By the normal form of $g\in\tilde{\Gamma}$, we mean the expression
$g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots z^{\epsilon_{l}}a_{l+1}$
with $a_{i}\in\Gamma$ and $\epsilon_{i}\in\{\pm\}$, which is reduced in the sense that if $i>1$ and $a_{i}=1$, then $\epsilon_{i-1}=\epsilon_{i}$.
We introduce the tree $\mathrm{T}\coloneq\tilde{\Gamma}/\Gamma$ on which $\tilde{\Gamma}$ acts. We declare that the points $s,t\in\mathrm{T}$ are adjacent if $s^{-1}t\in\{\Gamma z^{\pm}\Gamma\}$. With this graph structure $\mathrm{T}$ becomes a tree (of infinite degree). We call the point $o\coloneq\Gamma\in\mathrm{T}$ the origin. Every $t\in\mathrm{T}$ has unique representative $\rho(t)\in\tilde{\Gamma}$ whose normal form ends by $z^{\pm}$, or $\rho(o)=1$ if $t=o$. Note that $\rho(at)=a\rho(t)$ for all $a\in\Gamma$ and $t\in\mathrm{T}\setminus\{o\}$ and that $\rho(z^{\pm}t)=z^{\pm}\rho(t)$ for all $t\in\mathrm{T}$.
Suppose first that $s$ and $t$ are adjacent. We label the oriented edge from $t$ to $s$ by $\ell(t,s)\coloneq\kappa(\rho(t)^{-1}s)\in\Gamma\{z_{\pm}\}$, where $\kappa(az^{\pm}\Gamma)\coloneq az_{\pm}$. Note that $t$ and $\ell(t,s)$ uniquely determines $s$. E.g., $\ell(az\Gamma,azbz\Gamma)=bz_{+}$ and $\ell(azbz\Gamma,az\Gamma)=z_{-}$ for every $a,b\in\Gamma$. The Figure 1 shows how $\mathrm{T}$ and $\ell$ look like. Next we extend the label on $(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$ by setting $\ell(t,s)\coloneq\ell(t,t_{1})$, where $t_{1}$ is the unique point between $t$ and $s$ that is adjacent to $t$, or colloquially, $\ell(t,s)$ is the first $\Gamma\{z^{\pm}\}$ component in the normal form $\rho(t)^{-1}\rho(s)$. For example, $\ell(az,azbzcz\Gamma)=bz_{+}$ and $\ell(cz\Gamma,azbz\Gamma)=z_{-}$ unless $a=c$.
Figure 1. The tree $\mathrm{T}$
The following is not hard to see.
Lemma 3.
Let $(t,s)\in(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$. For $a\in\Gamma$, one has $\ell(at,as)=\ell(t,s)$ for $t\neq o$ and $\ell(o,as)=a\ell(o,s)$. Also, $\ell(zt,zs)=\ell(t,s)$.
Let $\Gamma\curvearrowright X$ be a topologically-free extreme boundary with an axial sequence $(z_{n})_{n}$. We denote by $\varphi_{n}\colon\tilde{\Gamma}\to\Gamma$ the homomorphism that is identity on $\Gamma$ and send $z$ to $z_{n}\in\Gamma$.
Lemma 4.
Let $g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots z^{\epsilon_{l}}$ be an element in the normal form with $l\geq 1$ and $x\in X\setminus\{z_{\bar{\epsilon}_{l}}\}$. Then $\varphi_{n}(g)x\to a_{1}z_{\epsilon_{1}}$ as $n\to\infty$.
Proof.
Let $U_{\pm}$ be arbitrarily small neighborhoods of $z_{\pm}$ that satisfy $x\notin U_{\bar{\epsilon}_{l}}$ and $a_{i}(U_{+}\cup U_{-})\cap(U_{+}\cup U_{-})=\emptyset$. Then by the ping-pong argument, one has $\varphi_{n}(g)x\in a_{1}U_{\epsilon_{1}}$ for all $n$ large enough. ∎
We denote by $\sigma$ the corresponding action of $\Gamma$ on $C(X)\subset C(X)^{\mathcal{U}}$, where the embedding is the constant embedding. We denote by $\sigma_{z}$ the automorphism arising from the sequence $(\sigma_{z_{n}})_{n}$. They together give rise to the action of $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$ on $C(X)^{\mathcal{U}}$, which is given by
$\sigma_{g}([f_{n}]_{n})=[\sigma_{\varphi_{n}(g)}(f_{n})]_{n}.$
We are interested in the commutative $\mathrm{C}^{*}$-algebra
$C(X_{\mathrm{T}})\coloneq\mathrm{C}^{*}(\bigcup_{g\in\tilde{\Gamma}}\sigma_{g}(C(X)))\subset C(X)^{\mathcal{U}}$
on which $\tilde{\Gamma}$ acts, where $X_{\mathrm{T}}$ is the Gelfand spectrum of $C(X_{\mathrm{T}})$. By definition, there is a surjective homomorphism
$Q\colon C(X^{\mathrm{T}})\cong\bigotimes_{t\in\mathrm{T}}C(X)\to C(X_{\mathrm{T}}),\quad Q(f^{(s)})=\sigma_{\rho(s)}(f),$
where $f^{(s)}\in C(X^{\mathrm{T}})$ is defined for $f\in C(X)$ and $\mathbf{x}=(x_{t})_{t}$ by $f^{(s)}(\mathbf{x})\coloneq f(x_{s})$. This gives rise to an embedding $Q_{*}\colon X_{T}\to X^{\mathrm{T}}$. We equip $X^{\mathrm{T}}$ a $\tilde{\Gamma}-$action by declaring that $a\in\Gamma$ acts at $\mathbf{x}=(x_{t})_{t}\in X^{\mathrm{T}}$ by $(a\mathbf{x})_{o}\coloneq ax_{o}$ and $(a\mathbf{x})_{t}\coloneq x_{a^{-1}t}$ for $t\in\mathrm{T}\setminus\{o\}$, and that $z$ acts by $(z\mathbf{x})_{t}\coloneq x_{z^{-1}t}$.
Lemma 5.
The embedding $Q_{*}\colon X_{\mathrm{T}}\to X^{\mathrm{T}}$ is $\tilde{\Gamma}-$equivariant.
Proof.
Let $\varsigma$ denote the corresponding $\tilde{\Gamma}-$action on $C(X^{\mathrm{T}})$. It is enough to check that
$Q(\varsigma_{a}(f^{(o)}))=Q(\sigma_{a}(f)^{(o)})=\sigma_{a}(f)=\sigma_{a}(Q(f^{(o)})),$
$Q(\varsigma_{a}(f^{(s)}))=Q(f^{(as)})=\sigma_{\rho(as)}(f)=\sigma_{a}(\sigma_{\rho(s)}(f))=\sigma_{a}(Q(f^{(s)}))$
for $s\neq o$ and,
$Q(\varsigma_{z}(f^{(s)}))=Q(f^{(zs)})=\sigma_{\rho(zs)}(f)=\sigma_{z}(\sigma_{\rho(s)}(f))=\sigma_{z}(Q(f^{(s)})).$
Hereafter, we omit writing $Q$ and identify $\varsigma$ with $\sigma$. Thus we view $X_{\mathrm{T}}$ as a compact subset of $X^{\mathrm{T}}$. For every $s\in\mathrm{T}$, we define the continuous embedding $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ by $\theta(s,x)_{s}\coloneq x$ and $\theta(s,x)_{t}\coloneq\ell(t,s)$ for $t\neq s$. We also define the continuous embedding $\theta\colon\partial\mathrm{T}\to X^{\mathrm{T}}$ by $\theta(\omega)_{t}\coloneq\ell(t,\omega)$. Note that $\omega\to\theta(\omega)_{t}$ is locally constant.
The following theorem asserts that $\Gamma\curvearrowright X_{\mathrm{T}}$ has a “tree structure”. See Figure 2.
Figure 2. The space $X_{\mathrm{T}}$ around $\theta(s,X)$, where $s=az\Gamma$ and $t=azbz\Gamma$, $\theta(s,bz_{+})=\theta(t,z_{-})$ and $\theta(s,z_{-})=\theta(o,az)$.
Theorem 6.
In the above setting for $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$, the following hold.
  1. (1)
    $X_{\mathrm{T}}=\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})$.
  2. (2)
    $\theta(\mathrm{T},X)\cap\theta(\partial\mathrm{T})=\emptyset$.
  3. (3)
    For $s\neq t$ and $x,y\in X$, one has $\theta(s,x)=\theta(t,y)$ if and only if $s$ and $t$ are adjacent and $x=\ell(s,t)$ and $y=\ell(t,s)$.
  4. (4)
    For $a\in\Gamma$, one has $a\theta(o,x)=\theta(o,ax)$ and $a\theta(s,x)=\theta(as,x)$ for $s\neq o$.
  5. (5)
    $z\theta(s,x)=\theta(zs,x)$.
  6. (6)
    $\theta\colon\partial\Gamma\to X_{\mathrm{T}}$ is $\tilde{\Gamma}-$equivariant.
  7. (7)
    A sequence $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(s,x)$ in $X_{\mathrm{T}}$ if and only if either $(t_{n})_{n}$ converges to $s$ in $\bar{\mathrm{T}}$ and the sequence $(y_{n})_{n}$, defined by $y_{n}=x_{n}$ if $t_{n}=s$ and $y_{n}=\ell(s,t_{n})$ if $t_{n}\neq s$, converges to $x$ in $X$; or there is an adjacent point $s_{1}$ to $s_{0}\coloneq s$ in $\mathrm{T}$ such that $t_{n}\in\{s_{0},s_{1}\}$ eventually, $x=\ell(s_{0},s_{1})$, and the subsequences $(x_{n})_{\{n:t_{n}=s_{i}\}}$ (possibly there is only one of them) converge to $\ell(s_{i},s_{j})$ for $\{i,j\}=\{0,1\}$. A sequence $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(\omega)$ in $X_{\mathrm{T}}$ if and only if $(t_{n})_{n}$ converges to $\omega$ in $\bar{\mathrm{T}}$.
  8. (8)
    A sequence $\theta(\omega_{n})$ converges to $\theta(s,x)$ in $X_{\mathrm{T}}$ if and only if $(\omega_{n})_{n}$ converges to $s$ in $\bar{\mathrm{T}}$ and and $(\ell(s,\omega_{n}))_{n}$ converges to $x$ in $X$.
  9. (9)
    One has $\theta(\mathrm{T},\Gamma\{z_{\pm}\})=\tilde{\Gamma}\theta(o,\{z_{\pm}\})$. The map $\beta\colon X_{\mathrm{T}}\setminus\tilde{\Gamma}\theta(o,\{z_{\pm}\})\to\bar{\mathrm{T}}$, given by $\beta(\theta(s,x))=s$ and $\beta(\theta(\omega))=\omega$, is $\tilde{\Gamma}-$equivariant and continuous.
Proof.
[…]
The assertions (4-6) follow from Lemma 3; E.g., for $a\in\Gamma$ and $s\in\mathrm{T}\setminus\{o\}$, one has
$(a\theta(s,x))_{o}=a\theta(s,x)_{o}=a\ell(o,s)=\ell(o,as)=\theta(as,x)_{o}$
and for $t\neq o$
$(a\theta(s,x))_{t}=\theta(s,x)_{a^{-1}t}=\left\{\begin{array}[]{ll}x&\mbox{ if $t=as$}\\ \ell(a^{-1}t,s)=\ell(t,as)&\mbox{ if $t\neq as$}\end{array}\right\}=\theta(as,x)_{t}.$
The “if” part of (7-8) are easy. For the “only if” part, suppose that $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(s,x)$. If $(t_{n})_{n}$ is converges to some point $s^{\prime}$, then by the “if” part and the compactness of $X$, one has $\theta(t_{n},x_{n})\to\theta(s^{\prime},y)$, where $y$ is the limit of $y_{n}\coloneq x_{n}$ if $t_{n}=s^{\prime}$ and $y_{n}\coloneq\ell(s^{\prime},t_{n})$ if $t_{n}\neq s^{\prime}$. Since $\theta(s^{\prime},y)=\theta(s,x)$, one has either $s^{\prime}=s$ or $s^{\prime}$ is an adjacent point to $s$ that satisfies $x=\ell(s,s^{\prime})$ and $y=\ell(s^{\prime},s)$. This proves the “only if” part for the case $(t_{n})_{n}$ is convergent. Now consider the case where $(t_{n})_{n}$ is not convergent. By the previous case, the only limit points are either $s$ or the unique adjacent point $s^{\prime}$ that satisfies $x=\ell(s,s^{\prime})$. The other claimed conditions also follow. The rest of the proof of (7-8) is similar.
We prove (1). We first prove that $\theta(\mathrm{T},X)\subset X_{\mathrm{T}}$. Let $s\in\mathrm{T}$ and $x\in X\setminus\Gamma\{z_{\pm}\}$ be given. We consider the character $\chi$ on $C(X_{\mathrm{T}})\subset C(X)^{\mathcal{U}}$ arising from the sequence $(\varphi_{n}(\rho(s))x)_{n}$. Then by Lemma 4 one has
$\chi(f^{(t)})=\lim\nolimits_{\mathcal{U}}\sigma_{\varphi_{n}(\rho(t))}(f)(\varphi_{n}(\rho(s))x)=\lim\nolimits_{\mathcal{U}}f(\varphi_{n}(\rho(t)^{-1}\rho(s))x)=f(\ell(t,s))$
for $t\neq s$ and $\chi(f^{(s)})=f(x)$. This means $\chi$ corresponds to the point $\theta(s,x)$ in $X_{\mathrm{T}}$. Since $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ is continuous and $X\setminus\Gamma\{z_{\pm}\}$ is dense in $X$, one sees that $\theta(s,X)\subset X_{\mathrm{T}}$. Similarly for $\theta(\partial\mathrm{T})$. It follows that $\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})\subset X_{\mathrm{T}}$.
It remains to prove the opposite inclusion. Let $\mathbf{x}\coloneq(x_{t})_{t}\in X^{\mathrm{T}}$ be given and assume that $\mathbf{x}\in X_{\mathrm{T}}$. We define $w\colon\mathrm{T}\to\mathrm{T}$, “the orientation of $\mathbf{x}$ at tt”, as follows. If $x_{t}\in X\setminus\Gamma\{z_{\pm}\}$, then $w(t)\coloneq t$. If $x_{t}\in\Gamma\{z_{\pm}\}$, then $w(t)\coloneq\rho(t)\kappa^{-1}(x_{t})$, where $\kappa^{-1}(az_{\pm})\coloneq az^{\pm}\Gamma\in\mathrm{T}$. Note that $w(t)$ is adjacent to $t$ in the latter case. We claim that if $s,t\in\mathrm{T}$ are adjacent, then it cannot happen that $w(s)\neq t$ and $w(t)\neq s$ simultaneously. Suppose for a contradiction that it happened. We may assume that $s$ is between $o$ and $t$ and $\rho(t)=\rho(s)az$ with no cancellation. Thus one has $x_{s}\neq az_{+}$ and $x_{t}\neq z_{-}$. Let $f\in C(X)$ be such that $f(x_{s})=1$ and $f=0$ on a neighborhood of $az_{+}$. Similarly let $g$ be such that $g(x_{t})=1$ and $g=0$ on a neighborhood of $z_{-}$. We consider the function $f^{(s)}g^{(t)}$ on $X^{\mathrm{T}}$. It satisfies $(f^{(s)}g^{(t)})(\mathbf{x})=f(x_{s})g(x_{t})=1$. However,
$(f^{(s)}g^{(t)})|_{X_{\mathrm{T}}}=Q(f^{(s)}g^{(t)})=\sigma_{\rho(s)}(f)\sigma_{\rho(t)}(g)=\sigma_{\rho(s)}(f\sigma_{az}(g))=0$
since $(\operatorname{supp}f)\cap az_{n}(\operatorname{supp}g)=\emptyset$ eventually. This means that $\mathbf{x}\in X^{\mathrm{T}}\setminus X_{\mathrm{T}}$, a contradiction. Thus we have proved the claim for adjacent points $s,t\in\mathrm{T}$ that $w(s)\neq t$ implies $w(t)=s$. This yields three possibilities. (1): There is $s\in\mathrm{T}$ such that $w$ is oriented to the $s$ direction. (2): There is $\omega\in\partial\mathrm{T}$ such that $w$ is oriented to the $\omega$ direction. (3): There is a pair $s_{1},s_{2}$ of adjacent points such that $w$ is oriented to $\{s_{1},s_{2}\}$ direction and $w(s_{i})=s_{j}$ for $\{i,j\}=\{1,2\}$. In the first case, one has $\mathbf{x}=\theta(s,x_{s})$. In the second case, one has $\mathbf{x}=\theta(\omega)$. In the third case, one has $\mathbf{x}=\theta(s_{1},\ell(s_{1},s_{2}))=\theta(s_{2},\ell(s_{2},s_{1}))$. It follows that $\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})\supset X_{\mathrm{T}}$.
The assertion (9) is a consequence of (1-8). ∎
4. Proof of Theorem 1
Proof of Theorem 1.
By Theorem 2.6 in [Ro], $\mathrm{C}^{*}$-selflessness follows from the “more precisely” statement. Let $\Gamma\curvearrowright X$ be a topologically-free extreme boundary with an axial sequence $(z_{n})_{n}$ and $\pi\colon\Gamma*\mathopen{\langle}z\mathclose{\rangle}\to\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ be the corresponding homomorphism. We set $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$. We take a $\Gamma$-equivariant embedding $C(X)\subset\ell_{\infty}\Gamma\subset\mathbb{B}(\ell_{2}\Gamma)$. It gives rise to $\tilde{\Gamma}-$equivariant embedding of $C(X_{\mathrm{T}})$ into $\mathbb{B}(\ell_{2}\Gamma)^{\mathcal{U}}$, where $C(X_{\mathrm{T}})$ is defined in the previous section. We take a faithful representation of $\mathbb{B}(\ell_{2}\Gamma)^{\mathcal{U}}$ on a Hilbert space $\mathcal{H}$ and view $\pi$ as a covariant representation of $\tilde{\Gamma}\ltimes C(X_{\mathrm{T}})$. Then the representation $\pi|_{C(X_{\mathrm{T}})}$ canonically extends to a normal representation $\pi$ of the $\mathrm{C}^{*}$-algebra $B(X_{\mathrm{T}})$ of the bounded Borel functions on $X_{\mathrm{T}}$. Let $p$ denote the orthogonal projection on $\mathcal{H}$ corresponding to the countable subset $\tilde{\Gamma}\{z_{\pm}\}$ of $X_{\mathrm{T}}$, which is possibly zero. The unitary representation $p\pi$ of $\tilde{\Gamma}$ is covariant with $\ell_{\infty}\tilde{\Gamma}$ and hence unitarily equivalent to a multiple of the regular representation. By Theorem 6, there is a $\tilde{\Gamma}-$equivariant embedding of $C(\bar{\mathrm{T}})$ into $p^{\perp}\pi(B(X_{\mathrm{T}}))$. Thus by Theorem 2 and the fact that $\pi|_{\Gamma}$ is weakly equivalent to the regular representation of $\Gamma$, the unitary representation $p^{\perp}\pi$ of $\tilde{\Gamma}$ is weakly contained in the regular representation. ∎
A crossed product analogue of Theorem 1 holds true under the additional assumption that $\Gamma$ is exact, which ensures $\Gamma\ltimes_{\min}A^{\mathcal{U}}\subset(\Gamma\ltimes_{\min}A)^{\mathcal{U}}$.