### v7

# Proximality and selflessness for group C*-algebras

Abstract.
We prove that the reduced group $\mathrm{C}^{*}$-algebras of infinite countable discrete groups having topologically-free extreme boundaries, or more generally groups that satisfy certain combinatorial property including all acylindrically hyperbolic groups with no nontrivial finite normal subgroups and all Zariski-dense subgroups of $\mathrm{PSL}(n,\mathbb{R})$, are selfless in the sense of L. Robert. This generalizes the recent results of Amrutam, Gao, Kunnawalkam Elayavalli, and Patchell, and of Vigdorovich. We also prove that selflessness is stable under tensor product among exact $\mathrm{C}^{*}$-algebras and that a $\mathrm{C}^{*}$-probability space is selfless provided that it is either simple and purely infinite or simple, exact, $\mathcal{Z}$-stable, and uniquely tracial.
Key words and phrases:
Extremely proximal actions, extreme boundaries, selfless $\mathrm{C}^{*}$-algebras
1991 Mathematics Subject Classification:
Primary 37B05; Secondary 46L35
The author was partially supported by JSPS KAKENHI Grant Numbers 24K00527, 25H00588, 25H00593
1. Introduction
A remarkable property called selflessness is recently introduced for a $\mathrm{C}^{*}$-algebra by L. Robert ([Ro]) and it quickly attracted a number of researchers’ attention (see, e.g., [AGKEP], [HKER], [RTV], [Vi]) as it implies many important regularity properties such as simplicity, (in the tracial setting) stable rank one, and strict comparison. In particular, Amrutam, Gao, Kunnawalkam Elayavalli, and Patchell ([AGKEP]) has proved that a large family of groups, namely acylindrically hyperbolic groups with the rapid decay property and with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless. Here, we say a group $\Gamma$ is $\mathrm{C}^{*}$-selfless if it its reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(\Gamma)$ is selfless.
In this paper, we prove that every infinite countable discrete group with a minimal topologically-free extremely-proximal action is $\mathrm{C}^{*}$-selfless. Our method is topological, as opposed to the more analytical one in [AGKEP], and to construct a kind of “tree-graded” space ([DS]) out of an extreme boundary (see Section 3). Note that every $\mathrm{C}^{*}$-selfless group is $\mathrm{C}^{*}$-simple ([Ro]) and whether the converse is also true is an open problem (Problem XCI in [STW]). We remind Kalantar and Kennedy’s theorem ([KK]) that a discrete group is $\mathrm{C}^{*}$-simple if and only if it has a minimal topological-free strongly-proximal action. As the names suggest, strong proximality is weaker than extreme proximality ([Gl]).
Definition.
Let $\Gamma$ be a countable discrete group and $\Gamma\curvearrowright X$ be an action on a compact topological space $X$. The action $\Gamma\curvearrowright X$ (or the $\Gamma$-space $X$) is called an extreme boundary if it is minimal and extremely proximal ([Gl]) in the sense that for every non-empty open subsets $U$ and $V$ of $X$, there is $g\in\Gamma$ such that $g(X\setminus U)\subset V$.
For $x\in X$, we denote by $\Gamma_{x}\coloneq\{g\in\Gamma:gx=x\}$ the stabilizer subgroup at $x$. The action $\Gamma\curvearrowright X$ is said to be topologically-free if the set of points with trivial stabilizer groups is dense in $X$. Note that if an extreme boundary $\Gamma\curvearrowright X$ is not topologically-free, then every proper closed subset of $X$ is pointwise fixed by some $g\in\Gamma\setminus\{1\}$.
We say a sequence $(z_{n})_{n}$ in $\Gamma$ is axial if there is a topologically-free extreme boundary $\Gamma\curvearrowright X$ with distinct points $z_{\pm}\in X$ that satisfies the following two conditions. (1): For every neighborhoods $U_{\pm}$ of $z_{\pm}$ one has $z_{n}(X\setminus U_{-})\subset U_{+}$ eventually, or equivalently that $z_{n}^{-1}(X\setminus U_{+})\subset U_{-}$ eventually. (2): The $\Gamma$-action on $\{z_{\pm}\}$ is free in the sense that $g\{z_{\pm}\}\cap\{z_{\pm}\}\neq\emptyset$ implies $g=1$. Note that a second countable extreme boundary $X$ with $|X|>2$ accommodates an axial sequence if and only if it is topologically-free.
Theorem 1.
An infinite countable discrete group $\Gamma$ having a topologically-free extreme boundary is $\mathrm{C}^{*}$-selfless. More precisely, for any axial sequence $(z_{n})_{n}$ in $\Gamma$ and any free ultrafilter $\mathcal{U}$ on $\mathbb{N}$, the homomorphism
$\Gamma*\mathopen{\langle}z\mathclose{\rangle}\to\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}},$
given by the diagonal embedding of $\Gamma$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$ and $z\mapsto[z_{n}]_{n}\in\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$, induces a faithful embedding of the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(G*\mathopen{\langle}z\mathclose{\rangle})$ into $\mathrm{C}^{*}_{\lambda}(\Gamma)^{\mathcal{U}}$.
See, e.g., [Bo], [BIO], [FLMMS], [IO], [JR], [LS], [LBMB] for examples of extreme boundaries. Every group satisfying the conclusion of Theorem 1 is mixed-identity-free (see [HO] for the terminology) and converse holds true if $\Gamma$ has a faithful minimal action of general type on a tree ([FLMMS]), since such an action gives rise to an extreme boundary. In particular, non-elementary free product groups are $\mathrm{C}^{*}$-selfless. For the case of amalgamated free products and HNN extensions, see [FLMMS] and [IO] for a characterization of topological freeness of the corresponding Bass–Serre tree compactifications. Non-elementarily relatively hyperbolic groups with no nontrivial finite normal subgroups are $\mathrm{C}^{*}$-selfless, as their Bowditch compactifications ([Bo]) are topologically-free extreme boundaries. It is not clear whether this generalizes to acylindrically hyperbolic groups (with no assumption on the rapid decay property, cf. [AGKEP]), because their hyperbolic structure need not yield suitable compactifications. We fix this problem in Section 8 and prove that a group satisfying property $\mathrm{P}_{\mathrm{PHP}}$ (defined there) is $\mathrm{C}^{*}$-selfless. The class of groups with property $\mathrm{P}_{\mathrm{PHP}}$ contains all acylindrically hyperbolic groups with no nontrivial finite normal subgroups as well as all Zariski-dense subgroups of $\mathrm{PSL}(d\geq 2,\mathbb{R})$ (cf. [Vi]).
There are $\mathrm{C}^{*}$-selfless groups outside the mixed-identity-free realm.
Theorem 2.
Let $(A_{i},\varphi_{i})$ be separable $\mathrm{C}^{*}$-probability spaces. Assume that all $(A_{i},\varphi_{i})$ are selfless and exact. Then the tensor product $\bigotimes_{i}(A_{i},\varphi_{i})$ is selfless. In particular, the class of countable discrete groups that are $\mathrm{C}^{*}$-selfless and exact is closed under direct product.
We introduce the notion of complete selflessness in Section 6 and prove that the examples in Theorem 1 and their tensor product are in fact completely selfless, while no nuclear tracial $\mathrm{C}^{*}$-probability space is completely selfless. The following is a partial converse to Theorem 3.1 in [Ro] and solves Question 5.4 in [Ro]. Note that by Matui and Sato’s theorem ([MS]), $\mathcal{Z}$-stability is necessary for a nuclear tracial $\mathrm{C}^{*}$-probability space to be selfless.
Theorem 3.
A simple and purely infinite $\mathrm{C}^{*}$-probability space $(A,\varphi)$ is completely selfless. A simple, exact, $\mathcal{Z}$-stable, and uniquely tracial $\mathrm{C}^{*}$-probability space $(A,\tau)$ is selfless.
Notations
We use the symbol $\pm$” for a slightly abusive notation for “++ and/or -” or “+1+1 and/or 1-1”. We denote by $A^{\mathcal{U}}$ the ultrapower of a $\mathrm{C}^{*}$-algebra $A$ and by $a_{n}]_{n}$ the element in $A^{\mathcal{U}}$ represented by a sequence $(a_{n})_{n}$ in $A$. A $\mathrm{C}^{*}$-probability space is a pair $(A,\varphi)$ of a unital $\mathrm{C}^{*}$-algebra $A$ and a distinguished state $\varphi$ which is not necessarily faithful but assumed GNS-faithful (a.k.a. non-degenerate) unless the $\mathrm{C}^{*}$-algebra in consideration is an ultrapower or an ultraproduct.
Acknowledgments
A part of this work was carried out while the author was visiting the Mathematisches Forschungsinstitut Oberwolfach for the workshop “C*-Algebras”, August 3–8, 2025 and the Isaac Newton Institute for the program “Operators, Graphs, Groups” in October 2025. He acknowledges the kind hospitality and the exciting environment provided by the institutes. He thanks G. Patchell for the stimulating talk at MFO and a useful comment on this paper, E. Breuillard and I. Vigdorovich for insights into $\mathrm{PSL}(d,\mathbb{R})$, and L. Robert for a comment.
2. Trees and covariant representations
We recall the notion of trees and their compactification. A tree is a connected graph without nontrivial cycles. See Section 5.2 and Appendix E in [BO] for more on this theme. Here we consider a countable tree $\mathrm{T}$ (of infinite degree). The tree $\mathrm{T}$ is identified with its vertex set. A geodesic ray is an infinite sequence $\omega(0),\omega(1),\ldots$ in $\mathrm{T}$ such that $\omega(n)$ adjacent to $\omega(n-1)$ and $\omega(n)\neq\omega(n-2)$. Geodesic rays $\omega$ and $\omega^{\prime}$ are equivalent if there are $N$ and $N^{\prime}$ such that $\omega(N+n)=\omega^{\prime}(N^{\prime}+n)$ for all $n\geq 0$. The boundary $\partial\mathrm{T}$ is the equivalence classes of geodesic rays and $\bar{\mathrm{T}}\coloneq\mathrm{T}\cup\partial\mathrm{T}$ is the compactification of $\mathrm{T}$ (whose topology will be introduced shortly). We define $f\colon\mathrm{T}\times\bar{\mathrm{T}}\to\mathrm{T}$ as follows. For $s\in\mathrm{T}$, we set $f(s,s)=s$. For distinct $s,t\in\mathrm{T}$, we define $f(t,s)\in\mathrm{T}$ to be the unique point between $t$ and $s$ that is adjacent to $t$. For $t\in\mathrm{T}$ and $\omega\in\partial\mathrm{T}$, we define $f(t,\omega)\coloneq f(t,\omega(n))$ for any $n$ large enough. For each $t\in\mathrm{T}$, we consider a copy $\mathrm{T}_{t}$ of $\mathrm{T}$ (or the union of $t$ and its adjacent points) and equip $\mathrm{T}_{t}$ a compact topology by taking the one-point compactification $\mathrm{T}\cup\{\infty\}$ of $\mathrm{T}$ and identify the new point $\infty$ with $t$. We consider the embedding
$\bar{\mathrm{T}}\ni\omega\mapsto(f(t,\omega))_{t}\in\prod_{t\in\mathrm{T}}\mathrm{T}_{t}$
and equip $\bar{\mathrm{T}}$ with the induced topology. This makes $\bar{\mathrm{T}}$ a second-countable compact topological space such that every automorphism on $\mathrm{T}$ extends to a homeomorphism on $\bar{\mathrm{T}}$ (Proposition 5.2.5 in [BO]). Note that if $s\in\mathrm{T}$ and $t_{n}\in\mathrm{T}$ are such that $f(s,t_{n})\to\infty$ in $\mathrm{T}$, then $t_{n}\to s$ in $\bar{\mathrm{T}}$.
Theorem 4.
Let $\Lambda$ be a countable discrete group acting on a countable tree $\mathrm{T}$ and on a $\mathrm{C}^{*}$-algebra $C$. Let $\pi\colon\Lambda\ltimes_{\max}C\to\mathbb{B}(\mathcal{H})$ be a covariant representation and assume that there is a $\Lambda$-equivariant representation of $C(\bar{\mathrm{T}})$ into the commutant $\pi(C)^{\prime}$ of $\pi(C)$. Then $\pi$ is continuous on the reduced crossed product $\Lambda\ltimes_{\min}C$ if for every $t]\in\mathrm{T}/\Lambda$ the restriction of $\pi$ to the stabilizer subgroup $\Lambda_{t}\coloneq\{g\in\Lambda:gt=t\}$ is continuous on the reduced crossed product $\Lambda_{t}\ltimes_{\min}C$.
Proof.
Let $\pi\colon\Lambda\ltimes_{\max}C\to\mathbb{B}(\mathcal{H})$ be given. We denote by $\tau$ the unitary representation of $\Lambda$ on $\ell_{2}\mathrm{T}$ induced by $\Lambda\curvearrowright\mathrm{T}$ and denote by $\sigma$ the $\Lambda$ action on $G$. We claim that $\pi$ is weakly contained in the representation $\tilde{\pi}$ on $\ell_{2}\mathrm{T}\otimes\mathcal{H}$ given by $g\mapsto(\tau\otimes\pi)(g)$ for $g\in\Lambda$ and $a\mapsto 1\otimes\pi(a)$ for $a\in C$. Indeed, there is a sequence $\zeta_{n}\colon\bar{\mathrm{T}}\to\mathrm{Prob}(\mathrm{T})$ of Borel maps such that $\sup_{x\in\bar{\mathrm{T}}}\|g\zeta_{n}^{x}-\zeta_{n}^{gx}\|_{1}\to 0$ for every $g\in\Lambda$ (Lemma 5.2.6 in [BO]). It follows that $h_{n}^{t}(x)\coloneq\zeta_{n}^{x}(t)$ defines sequences of positive Borel functions on $\bar{\mathrm{T}}$ such that $\sum_{t\in\mathrm{T}}h_{n}^{t}=1$ and
$\|\sum_{t\in\mathrm{T}}|h_{n}^{gt}-\sigma_{g}(h_{n}^{t})|\|_{\infty}\to 0.$
NB: the above sum is an infinite sum of positive Borel functions. The $\Lambda$-equivariant representation of $C(\bar{\mathrm{T}})$ into $\pi(C)^{\prime}$ uniquely extends to a $\sigma$-normal $\Lambda$-equivariant representation $\theta$ from the $\mathrm{C}^{*}$-algebra $B(\bar{\mathrm{T}})$ of bounded Borel functions into $\pi(C)^{\prime}$. By $\sigma$-normal”, we mean that it sends a bounded pointwise convergent sequence to a strong operator topology convergent sequence. We define isometries $V_{n}\colon\mathcal{H}\to\ell_{2}\mathrm{T}\otimes\mathcal{H}$ by $V_{n}\xi\coloneq\sum_{t\in\mathrm{T}}\delta_{t}\otimes\theta(h_{n}^{t})^{1/2}\xi$, or equivalently, $V_{n}$ is the column operator $(\theta(h_{n}^{t})^{1/2})_{t\in\mathrm{T}}$. Then
$\|V_{n}\pi(g)-(\tau\otimes\pi)(g)V_{n}\|^{2}=\|\sum_{t}|\theta(h_{n}^{gt})^{1/2}-\theta(\sigma_{g}(h_{n}^{t}))^{1/2}|^{2}\|_{\infty}\to 0$
for every $g\in\Lambda$ and $V_{n}\pi(a)=(1\otimes\pi(a))V_{n}$ for $a\in C$. This proves the claim that $\pi$ is weakly contained in $\tilde{\pi}$. Thus continuity of $\pi$ follows from that of $\tilde{\pi}$.
For the continuity of $\tilde{\pi}$, it suffices to consider the continuity of the vector states $\varphi$ associated with the vectors of the form $\delta_{t}\otimes\xi$, since they form a cyclic family. But $\varphi=\psi\circ E$, where $E$ is the canonical conditional expectation from $\Lambda\ltimes_{\max}C$ onto $\Lambda_{t}\ltimes_{\max}C$ and $\psi$ is the vector state associated with $\xi$. (We do not need the fact that the natural embedding $\Lambda_{t}\ltimes_{\max}C\subset\Lambda\ltimes_{\max}C$ is faithful.) Hence the continuity on the reduced crossed product $\Lambda\ltimes_{\min}C$ follows from that for $\Lambda_{t}\ltimes_{\min}C$ for $t\in\mathrm{T}$ one from each $\Lambda$-orbit. ∎
3. Extreme boundaries and “tree-graded” spaces
Let $\tilde{\Gamma}\coloneq\Gamma*\mathopen{\langle}z\mathclose{\rangle}$. By the reduced form of $g\in\tilde{\Gamma}$, we mean the unique expression
$g=a_{1}z^{\epsilon_{1}}a_{2}z^{\epsilon_{2}}\cdots a_{l}z^{\epsilon_{l}}a_{l+1}$
with $a_{i}\in\Gamma$ and $\epsilon_{i}\in\{\pm\}$ such that $i>1$ and $a_{i}=1$ implies $\epsilon_{i-1}=\epsilon_{i}$. We introduce the tree structure on $\mathrm{T}\coloneq\tilde{\Gamma}/\Gamma$ on which $\tilde{\Gamma}$ acts. We declare that the points $s,t\in\mathrm{T}$ are adjacent if $s^{-1}t\in\{\Gamma z^{\pm}\Gamma\}$. With this graph structure $\mathrm{T}$ becomes a tree (of infinite degree). We call the point $o\coloneq\Gamma\in\mathrm{T}$ the origin. Every $t\in\mathrm{T}$ has unique representative $\rho(t)\in\tilde{\Gamma}$ whose reduced form ends by $z^{\pm}$, or $\rho(o)=1$ if $t=o$. Note that $\rho(at)=a\rho(t)$ for all $a\in\Gamma$ and $t\in\mathrm{T}\setminus\{o\}$ and that $\rho(z^{\pm}t)=z^{\pm}\rho(t)$ for all $t\in\mathrm{T}$.
Suppose first that $s$ and $t$ are adjacent. We label the oriented edge from $t$ to $s$ by $\ell(t,s)\coloneq\kappa(\rho(t)^{-1}s)\in\Gamma\{z^{\pm}\}$, where $\kappa(az^{\pm}\Gamma)\coloneq az^{\pm}$. Note that $t$ and $\ell(t,s)$ together uniquely determines $s$. E.g., $\ell(az\Gamma,azbz\Gamma)=bz$ and $\ell(azbz\Gamma,az\Gamma)=z^{-1}$ for every $a,b\in\Gamma$. The Figure 1 shows how $\mathrm{T}$ and $\ell$ look like. Next we extend the label map on $(\mathrm{T}\times\bar{\mathrm{T}})\setminus\{\mbox{diagonal}\}$ by setting $\ell(t,s)\coloneq\ell(t,t_{1})$, where $t_{1}$ is the unique point between $t$ and $s$ that is adjacent to $t$, or colloquially, $\ell(t,s)$ is the first $\Gamma\{z^{\pm}\}$ component in the reduced form of $\rho(t)^{-1}\rho(s)$. For example, $\ell(az,azbzcz\Gamma)=bz$ and $\ell(cz\Gamma,azbz\Gamma)=z^{-1}$ unless $a=c$.
Figure 1. The tree $\mathrm{T}$
The following is not hard to see.
Lemma 5.
Let $(t,s)\in(\mathrm{T}\times\bar{\mathrm{T}})$ be a non-diagonal point. For $a\in\Gamma$, one has $\ell(o,as)=a\ell(o,s)$ and $\ell(at,as)=\ell(t,s)$ for $t\neq o$. Also, $\ell(zt,zs)=\ell(t,s)$.
Let $\Gamma\curvearrowright X$ be an extreme boundary with an axial sequence $(z_{n})_{n}$. We denote by $\varphi_{n}\colon\tilde{\Gamma}\to\Gamma$ the homomorphism that is identity on $\Gamma$ and sends $z$ to $z_{n}\in\Gamma$. A simple ping-pong argument on $X$ shows that $(\varphi_{n})_{n}$ is asymptotically injective and hence $\Gamma$ is mixed-identity-free. More precisely, the following is true.
Lemma 6.
Let $g=a_{1}z^{\epsilon_{1}}\cdots z^{\epsilon_{l}}a_{l+1}$ be an element in the reduced form with $l\geq 1$ and $x\in X\setminus\{a_{l+1}^{-1}z_{\bar{\epsilon}_{l}}\}$, where $\bar{\epsilon}$ is the opposite of $\epsilon$. Then $\varphi_{n}(g)x\to a_{1}z_{\epsilon_{1}}$ as $n\to\infty$.
We denote by $\sigma$ the corresponding action of $\Gamma$ on $C(X)\subset C(X)^{\mathcal{U}}$, where the embedding is the diagonal embedding. We denote by $\sigma_{z}$ the automorphism arising from the sequence $(\sigma_{z_{n}})_{n}$. They together give rise to the action of $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}$ on $C(X)^{\mathcal{U}}$, which is given by $\sigma_{g}([f_{n}]_{n})=[\sigma_{\varphi_{n}(g)}(f_{n})]_{n}$. We are interested in the $\tilde{\Gamma}-$invariant $\mathrm{C}^{*}$-subalgebra
$C(X_{\mathrm{T}})\coloneq\mathrm{C}^{*}(\bigcup_{g\in\tilde{\Gamma}}\sigma_{g}(C(X)))\subset C(X)^{\mathcal{U}},$
where $X_{\mathrm{T}}$ is the Gelfand spectrum of $C(X_{\mathrm{T}})$. By definition, there is a surjective homomorphism
$Q\colon C(X^{\mathrm{T}})\cong\bigotimes_{t\in\mathrm{T}}C(X)\to C(X_{\mathrm{T}}),\quad Q(f^{(s)})=\sigma_{\rho(s)}(f),$
where $f^{(s)}\in C(X^{\mathrm{T}})$ is defined for $f\in C(X)$ by $f^{(s)}(\mathbf{x})\coloneq f(x_{s})$, $\mathbf{x}=(x_{t})_{t}$. This gives rise to an embedding $Q_{*}\colon X_{T}\to X^{\mathrm{T}}$. We equip $X^{\mathrm{T}}$ a $\tilde{\Gamma}-$action by declaring that $a\in\Gamma$ acts at $\mathbf{x}=(x_{t})_{t}\in X^{\mathrm{T}}$ by $(a\mathbf{x})_{o}\coloneq ax_{o}$ and $(a\mathbf{x})_{t}\coloneq x_{a^{-1}t}$ for $t\in\mathrm{T}\setminus\{o\}$, and that $z$ acts by $(z\mathbf{x})_{t}\coloneq x_{z^{-1}t}$.
Lemma 7.
The embedding $Q_{*}\colon X_{\mathrm{T}}\to X^{\mathrm{T}}$ is $\tilde{\Gamma}-$equivariant.
Proof.
Let $\varsigma$ denote the corresponding $\tilde{\Gamma}-$action on $C(X^{\mathrm{T}})$. It is enough to check that
$Q(\varsigma_{a}(f^{(o)}))=Q(\sigma_{a}(f)^{(o)})=\sigma_{a}(f)=\sigma_{a}(Q(f^{(o)})),$
$Q(\varsigma_{a}(f^{(s)}))=Q(f^{(as)})=\sigma_{\rho(as)}(f)=\sigma_{a}(\sigma_{\rho(s)}(f))=\sigma_{a}(Q(f^{(s)}))$
for $s\neq o$ and,
$Q(\varsigma_{z}(f^{(s)}))=Q(f^{(zs)})=\sigma_{\rho(zs)}(f)=\sigma_{z}(\sigma_{\rho(s)}(f))=\sigma_{z}(Q(f^{(s)})).$
Hereafter, we omit writing $Q$ and identify $\varsigma$ with $\sigma$. Thus we view $X_{\mathrm{T}}$ as a compact subset of $X^{\mathrm{T}}$. We view the label map $\ell$ as a map taking values in $X$ by identifying $\Gamma\{z^{\pm}\}$ with $\Gamma\{z_{\pm}\}$. For each $s\in\mathrm{T}$, we define the continuous embedding $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ by $\theta(s,x)_{s}\coloneq x$ and $\theta(s,x)_{t}\coloneq\ell(t,s)$ for $t\neq s$. We also define the continuous embedding $\theta\colon\partial\mathrm{T}\to X^{\mathrm{T}}$ by $\theta(\omega)_{t}\coloneq\ell(t,\omega)$. Note that $\omega\to\theta(\omega)_{t}$ is locally constant for every $t$.
The space $X_{\mathrm{T}}$ has a kind of “tree-graded” structure ([DS]), as follows. See Figure 2.
Figure 2. The space $X_{\mathrm{T}}$ around $\theta(az\Gamma,bz_{+})=\theta(azbz\Gamma,z_{-})$.
Theorem 8.
In the above setting for $\tilde{\Gamma}=\Gamma*\mathopen{\langle}z\mathclose{\rangle}\curvearrowright\mathrm{T}$, the following hold.
  1. (1)
    $X_{\mathrm{T}}=\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})$.
  2. (2)
    $\theta(\mathrm{T},X)\cap\theta(\partial\mathrm{T})=\emptyset$.
  3. (3)
    For $s\neq t$ and $x,y\in X$, one has $\theta(s,x)=\theta(t,y)$ if and only if $s$ and $t$ are adjacent and $x=\ell(s,t)$ and $y=\ell(t,s)$.
  4. (4)
    For $a\in\Gamma$, one has $a\theta(o,x)=\theta(o,ax)$ and $a\theta(s,x)=\theta(as,x)$ for $s\neq o$.
  5. (5)
    $z\theta(s,x)=\theta(zs,x)$.
  6. (6)
    $\theta\colon\partial\Gamma\to X_{\mathrm{T}}$ is $\tilde{\Gamma}-$equivariant.
  7. (7)
    A sequence $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(s,x)$ in $X_{\mathrm{T}}$ if and only if either $(t_{n})_{n}$ converges to $s$ in $\bar{\mathrm{T}}$ and the sequence $(y_{n})_{n}$, defined by $y_{n}=x_{n}$ if $t_{n}=s$ and $y_{n}=\ell(s,t_{n})$ if $t_{n}\neq s$, converges to $x$ in $X$; or there is an adjacent point $s_{1}$ to $s_{0}\coloneq s$ in $\mathrm{T}$ such that $t_{n}\in\{s_{0},s_{1}\}$ eventually, $x=\ell(s_{0},s_{1})$, and the subsequences $(x_{n})_{\{n:t_{n}=s_{i}\}}$ (possibly there is only one of them) converge to $\ell(s_{i},s_{j})$ for $\{i,j\}=\{0,1\}$. A sequence $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(\omega)$ in $X_{\mathrm{T}}$ if and only if $(t_{n})_{n}$ converges to $\omega$ in $\bar{\mathrm{T}}$.
  8. (8)
    A sequence $\theta(\omega_{n})$ converges to $\theta(s,x)$ in $X_{\mathrm{T}}$ if and only if $(\omega_{n})_{n}$ converges to $s$ in $\bar{\mathrm{T}}$ and and $(\ell(s,\omega_{n}))_{n}$ converges to $x$ in $X$.
  9. (9)
    One has $\theta(\mathrm{T},\Gamma\{z_{\pm}\})=\tilde{\Gamma}\theta(o,\{z_{\pm}\})$. The map $\beta\colon X_{\mathrm{T}}\setminus\tilde{\Gamma}\theta(o,\{z_{\pm}\})\to\bar{\mathrm{T}}$, given by $\beta(\theta(s,x))=s$ and $\beta(\theta(\omega))=\omega$, is $\tilde{\Gamma}-$equivariant and continuous.
Proof.
[…]
The assertions (4-6) follow from Lemma 5; E.g., for $a\in\Gamma$ and $s\in\mathrm{T}\setminus\{o\}$, one has
$(a\theta(s,x))_{o}=a\theta(s,x)_{o}=a\ell(o,s)=\ell(o,as)=\theta(as,x)_{o}$
and for $t\neq o$
$(a\theta(s,x))_{t}=\theta(s,x)_{a^{-1}t}=\left\{\begin{array}[]{ll}x&\mbox{ if $t=as$}\\ \ell(a^{-1}t,s)=\ell(t,as)&\mbox{ if $t\neq as$}\end{array}\right\}=\theta(as,x)_{t}.$
The “if” part of (7-8) are easy. For the “only if” part, suppose that $(\theta(t_{n},x_{n}))_{n}$ converges to $\theta(s,x)$. If $(t_{n})_{n}$ is converges to some point $s^{\prime}$, then by the “if” part and the compactness of $X$, one has $\theta(t_{n},x_{n})\to\theta(s^{\prime},y)$, where $y$ is the limit of $y_{n}\coloneq x_{n}$ if $t_{n}=s^{\prime}$ and $y_{n}\coloneq\ell(s^{\prime},t_{n})$ if $t_{n}\neq s^{\prime}$. Since $\theta(s^{\prime},y)=\theta(s,x)$, one has by (3) either $s^{\prime}=s$ or $s^{\prime}$ is an adjacent point to $s$ that satisfies $x=\ell(s,s^{\prime})$ and $y=\ell(s^{\prime},s)$. This proves the “only if” part for the case $(t_{n})_{n}$ is convergent. Now consider the case where $(t_{n})_{n}$ is not convergent. By the previous case, the only limit points are either $s$ or the unique adjacent point $s^{\prime}$ that satisfies $x=\ell(s,s^{\prime})$. The other claimed conditions also follow. The rest of the proof of (7-8) is similar.
We prove (1). We first prove that $\theta(\mathrm{T},X)\subset X_{\mathrm{T}}$. Let $s\in\mathrm{T}$ and $x\in X\setminus\Gamma\{z_{\pm}\}$ be given. We consider the character $\chi$ on $C(X_{\mathrm{T}})\subset C(X)^{\mathcal{U}}$ arising from the sequence $(\varphi_{n}(\rho(s))x)_{n}$. Then by Lemma 6 one has
$\chi(f^{(t)})=\lim\nolimits_{\mathcal{U}}\sigma_{\varphi_{n}(\rho(t))}(f)(\varphi_{n}(\rho(s))x)=\lim\nolimits_{\mathcal{U}}f(\varphi_{n}(\rho(t)^{-1}\rho(s))x)=f(\ell(t,s))$
for $t\neq s$ and $\chi(f^{(s)})=f(x)$. This means $\chi$ corresponds to the point $\theta(s,x)$ in $X_{\mathrm{T}}$. Since $\theta(s,\,\cdot\,)\colon X\to X^{\mathrm{T}}$ is continuous and $X\setminus\Gamma\{z_{\pm}\}$ is dense in $X$, one sees that $\theta(s,X)\subset X_{\mathrm{T}}$. Similarly for $\theta(\partial\mathrm{T})$. It follows that $\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})\subset X_{\mathrm{T}}$.
It remains to prove the opposite inclusion. Let $\mathbf{x}\coloneq(x_{t})_{t}\in X^{\mathrm{T}}$ be given and assume that $\mathbf{x}\in X_{\mathrm{T}}$. We define $w\colon\mathrm{T}\to\mathrm{T}$, “the orientation of $\mathbf{x}$ at tt”, as follows. If $x_{t}\in X\setminus\Gamma\{z_{\pm}\}$, then $w(t)\coloneq t$. If $x_{t}\in\Gamma\{z_{\pm}\}$, then we define $w(t)$ to be the unique point that is adjacent to $t$ and satisfies $\ell(t,w(t))=x_{t}$. We claim that if $s,t\in\mathrm{T}$ are adjacent, then it cannot happen that $w(s)\neq t$ and $w(t)\neq s$ simultaneously. Suppose for a contradiction that it happened. We may assume that $s$ is between $o$ and $t$ and $\rho(t)=\rho(s)az$ with no cancellation. Thus one has $x_{s}\neq az_{+}$ and $x_{t}\neq z_{-}$. Let $f\in C(X)$ be such that $f(x_{s})=1$ and $f=0$ on a neighborhood of $az_{+}$. Similarly let $g$ be such that $g(x_{t})=1$ and $g=0$ on a neighborhood of $z_{-}$. We consider the function $f^{(s)}g^{(t)}$ on $X^{\mathrm{T}}$. It satisfies $(f^{(s)}g^{(t)})(\mathbf{x})=f(x_{s})g(x_{t})=1$. However,
$(f^{(s)}g^{(t)})|_{X_{\mathrm{T}}}=Q(f^{(s)}g^{(t)})=\sigma_{\rho(s)}(f)\sigma_{\rho(t)}(g)=\sigma_{\rho(s)}(f\sigma_{az}(g))=0$
since $(\operatorname{supp}f)\cap az_{n}(\operatorname{supp}g)=\emptyset$ eventually. This means that $\mathbf{x}\in X^{\mathrm{T}}\setminus X_{\mathrm{T}}$, a contradiction. Thus we have proved the claim for adjacent points $s,t\in\mathrm{T}$ that $w(s)\neq t$ implies $w(t)=s$. This yields three possibilities. (1): There is $s\in\mathrm{T}$ such that $w$ is oriented to the $s$ direction. (2): There is $\omega\in\partial\mathrm{T}$ such that $w$ is oriented to the $\omega$ direction. (3): There is a pair $s_{1},s_{2}$ of adjacent points such that $w$ is oriented to $\{s_{1},s_{2}\}$ direction and $w(s_{i})=s_{j}$ for $\{i,j\}=\{1,2\}$. In the first case, one has $\mathbf{x}=\theta(s,x_{s})$. In the second case, one has $\mathbf{x}=\theta(\omega)$. In the third case, one has $\mathbf{x}=\theta(s_{1},\ell(s_{1},s_{2}))=\theta(s_{2},\ell(s_{2},s_{1}))$. It follows that $\theta(\mathrm{T},X)\cup\theta(\partial\mathrm{T})\supset X_{\mathrm{T}}$.
Finally, the assertion (9) is a consequence of (1-8). ∎
4. Proof of Theorem 1
Proof of Theorem 1.
[…]
5. Direct products of $\mathrm{C}^{*}$-selfless groups
It is not hard to see that if $A\subset B$ is an existential embedding of $\mathrm{C}^{*}$-algebras (see Section 1 in [Ro]), then so is $A\otimes C\subset B\otimes C$, provided that the $\mathrm{C}^{*}$-algebra $C$ is exact. (Without the exactness assumption, the canonical embedding $A^{\mathcal{U}}\otimes C\subset(A\otimes C)^{\mathcal{U}}$ is not continuous w.r.t. the minimal tensor product. See [BO].) Hence, if $A_{i}\subset B_{i}$ are existential embedding and $B_{i}$ are exact, then the corresponding embedding $A_{1}\otimes A_{2}\subset B_{1}\otimes B_{2}$ is existential. The same holds true in the $\mathrm{C}^{*}$-probability space setting.
We denote by $(\mathcal{T},\omega)$ the Toeplitz $\mathrm{C}^{*}$-probability space consisting of the Toeplitz algebra generated by the unilateral shift $T$ on $\ell_{2}(\{0,1\,\ldots\})$ and the vacuum state $\omega$ associated with the vacuum vector $\delta_{0}$. We denote by $(\mathcal{C}_{1},\tau)\subset(\mathcal{T},\omega)$ the $\mathrm{C}^{*}$-probability space generated by $s\coloneq(T+T^{*})/2$ in $\mathcal{T}$. Note that $s$ corresponds to the identity function under the isomorphism $(\mathcal{C}_{1},\tau)\cong(C([-1,1]),\gamma)$, where the state $\gamma$ is given by by the semicircle probability measure $\frac{2}{\pi}\sqrt{1-s^{2}}\,\mathrm{d}s$. See Section 2.6 in [VDN].
We denote by $\mathbf{F}_{\infty}$ the free group on generators $\{t_{i}\}_{i=1}^{\infty}$. We view it as embedded in the reduced group $\mathrm{C}^{*}$-algebra $\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})$ acting on $\ell_{2}\mathbf{F}_{\infty}$. It forms the $\mathrm{C}^{*}$-probability space $(\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty}),\tau)$, where $\tau$ is the canonical tracial state that is given by $\tau(g)=\delta_{g,1}$ for $g\in\mathbf{F}_{\infty}$. It is well-known ([VDN]) that the sequence
$s_{n}=(8n)^{-1/2}\sum_{i=1}^{n}(t_{i}+t_{i}^{-1})$
in $(\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty}),\tau)$ strongly converges to $s$ in $(\mathcal{C}_{1},\tau)$. Namely, for every polynomial $p$, one has $\tau(p(s_{n}))\to\gamma(p(s))$ and $\|p(s_{n})\|\to\|p(s)\|$. We will prove a vectorial version of this fact in Theorem 9. Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space on which $\mathbf{F}_{\infty}$ acts (leaving $\varphi$ invariant). Then the reduced crossed product $A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty}$ is equipped with the canonical state, still denoted by $\varphi$, that is given by $\varphi(ag)=\varphi(a)\tau(g)$ for $a\in A$ and $g\in\mathbf{F}_{\infty}$.
Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space. We view $A$ as embedded in the ultrapower $(A,\varphi)^{\mathcal{U}}\coloneq(A^{\mathcal{U}},\varphi^{\mathcal{U}})$ diagonally. (The difference in the free ultrafilter $\mathcal{U}$ will have no effect on the discussion.) Consider a bounded sequence $(a_{n})_{n=1}^{\infty}$ of self-adjoint elements in $(A,\varphi)$ that is strongly convergent to $a$ in $(\mathrm{C}^{*}(\{1,a\}),\psi)$. We say the sequence $(a_{n})_{n=1}^{\infty}$ is strongly asymptotically free (w.r.t. $\mathcal{U}$) from a $\mathrm{C}^{*}$-subalgebra $A_{0}$ if the induced homomorphism
$(A_{0},\varphi|_{A_{0}})\ast(\mathrm{C}^{*}(\{1,a\}),\psi)\to(A^{\mathcal{U}},\varphi^{\mathcal{U}})$
is well-defined and continuous on the reduced free product. Note that the reduced free product is compatible with embedding (see [BD] or Corollary 4.8.4 in [BO].
Theorem 9.
Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space on which $\mathbf{F}_{\infty}$ acts by $\sigma$. Assume that
$\forall a\in A\quad\frac{1}{2n}\sum_{i=1}^{n}(\sigma_{t_{i}}(a)+\sigma_{t_{i}}^{-1}(a))\to\varphi(a).$
Then, the sequence $(s_{n})_{n=1}^{\infty}$ as above in $A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty}$ is strongly asymptotically free from $A$ (in fact from $A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty}$).
Proof.
The proof is inspired from [VDN] and [HP]. We can write $\mathbf{F}_{\infty}$ as a disjoint union
$\mathbf{F}_{\infty}=\{1\}\cup\bigcup_{i}(E_{i}^{+}\cup E_{i}^{-}),$
where $E_{i}^{\pm}$ is the set of those elements whose reduced form starts with (a positive power of) $t_{i}^{\pm}$. We denote by $P_{i,\pm}$ the orthogonal projection on $\ell_{2}\mathbf{F}_{\infty}$ that corresponds to $E_{i}^{\pm}$. They are mutually orthogonal. One has $t_{i}=P_{i,+}t_{i}+t_{i}P_{i,-}\eqcolon u_{i,+}+u_{i,-}^{*}$ (see Proof of Proposition 1.1 in [HP]). Note that $u_{i,\pm}$ are partial isometries such that $u_{i,\pm}^{*}u_{i,\pm}=1-P_{i,\mp}$ and $u_{i,\pm}u_{i,\pm}^{*}=P_{i,\pm}$. Hence, for
$T_{n}\coloneq(2n)^{-1/2}\sum_{i=1}^{n}(u_{i,+}+u_{i,-}),$
one has $s_{n}=(T_{n}+T_{n}^{*})/2$ and $T_{n}^{*}T_{n}=1-(2n)^{-1}\sum_{i=1}^{n}(P_{i,-}+P_{i,+})\approx 1$.
We represent $A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty}$ on the Hilbert space $\mathcal{H}\coloneq L^{2}(A,\varphi)\otimes\ell_{2}\mathbf{F}_{\infty}$ by $g\mapsto(1\otimes g)$ for $g\in\mathbf{F}_{\infty}$ and $a\mapsto\pi(a)$, $\pi(a)(\xi\otimes\delta_{h})=\sigma_{h}^{-1}(a)\xi\otimes\delta_{h}$, for $a\in A$. Then, by the assumption,
$(1\otimes T_{n})^{*}\pi(a)(1\otimes T_{n})=\pi(\frac{1}{2n}\sum_{i=1}^{n}(\sigma_{g_{i}}^{-1}(a)(1-P_{i,-})+\sigma_{g_{i}}(a)(1-P_{i,+}))\to\varphi(a)$
for $a\in A$. Hence $T\coloneq[(1\otimes T_{n})]_{n}\in\mathbb{B}(\mathcal{H})^{\mathcal{U}}$ is an isometry that satisfies $T^{*}\pi(a)T=\varphi(a)$ for $a\in A$. We denote by $\omega$ the vector state on $\mathbb{B}(\mathcal{H})$ associated with the vector $\hat{1}\otimes\delta_{1}$. Thus $(A,\varphi)\subset(\mathbb{B}(\mathcal{H}),\omega)$. The isometry $T$ in $(\mathbb{B}(\mathcal{H}),\omega)^{\mathcal{U}}$ generates a copy of the Toeplitz probability space $(\mathcal{T},\omega)$. In particular, $s_{n}]_{n}=(T+T^{*})/2$ generates a copy of $(\mathcal{C}_{1},\tau)$ in $(A\rtimes\mathbf{F}_{\infty},\varphi)^{\mathcal{U}}$. It is rather easy to see that $A$ and $\mathrm{C}^{*}(T)$ are free from each other w.r.t. $\omega^{\mathcal{U}}$. It remains to show continuity of the induced homomorphism from the reduced free product $(A,\varphi)\ast(\mathcal{T},\omega)$ into $(\mathbb{B}(\mathcal{H}),\omega)^{\mathcal{U}}$.
The universal $\mathrm{C}^{*}$-algebra generated by $A$ and an isometry $T$ that satisfy $T^{*}aT=\varphi(a)$ is the Toeplitz–Pimsner algebra $\mathcal{T}_{A,\varphi}$ over the Hilbert $A$-module $\mathop{\overline{\mathrm{span}}}ATA$ (see, e.g., Example 4.6.11 in [BO]). Since the canonical conditional expectation from $\mathcal{T}_{A,\varphi}$ onto $A$ is non-degenerate (Theorem 4.6.6 in [BO]), the Toeplitz–Pimsner algebra $\mathcal{T}_{A,\varphi}$ is isomorphic to the reduced free product $(A,\varphi)*(\mathcal{T},\omega)$. This proves the desired continuity. ∎
Proof of Theorem 2.
By Theorem 4.1 in [Ro], we may assume that $i\in\{1,2\}$. Since $(A_{i},\varphi_{i})$ are selfless, the embeddings
[…]
are existential (Theorem 2.6 in [Ro]). The $\mathrm{C}^{*}$-algebras $B_{i}$ are exact by Dykema’s theorem ([Dy] or Corollary 4.8.3 in [BO]). Hence the embedding of $(A_{1}\otimes A_{2},\varphi_{1}\otimes\varphi_{2})$ into $(B_{1}\otimes B_{2},\psi_{1}\otimes\psi_{2})$ is existential. We write $\mathbf{F}_{\infty}^{(i)}$ for the copy of $\mathbf{F}_{\infty}$ in $B_{i}$. Note that
[…]
$\supset(A\rtimes_{\mathrm{r}}\mathbf{F}_{\infty},\varphi),$
[…]
$\|\sum_{j=1}^{m}x_{j}\otimes y_{j}\|\leq 3(\sum_{j=1}^{m}\|x_{j}\|^{2}\|y_{j}\|^{2})^{1/2}\precsim m^{1/2}\max\|x_{j}\|\|y_{j}\|$
whenever $x_{j}$’s (or $y_{j}$’s) are freely independent and mean zero. ∎
6. Completely selfless $\mathrm{C}^{*}$-algebras
In this section, we indicate the way to circumvent the exactness assumption in Theorem 2. Recall that an embedding $A\subset B$ of $\mathrm{C}^{*}$-algebras is said to be existential if there is an ultrafilter $\mathcal{U}$ and an embedding $\sigma\colon B\hookrightarrow A^{\mathcal{U}}$ whose restriction to $A$ is the diagonal embedding of $A$ into $A^{\mathcal{U}}$. We say that the embedding is completely existential if there is $\sigma$ as above that moreover satisfies that for every $\mathrm{C}^{*}$-algebra $C$ the induced homomorphism $B\otimes C\to(A\otimes C)^{\mathcal{U}}$ is continuous. By the Effros–Haagerup lifting theorem, the additional condition is equivalent to that the homomorphism $\sigma$ is locally liftable to completely positive contractions into the $\ell_{\infty}$-sum $\prod A$. Namely, $A\subset B$ is completely existential if and only if for every finite dimensional operator system $E\subset B$ and $\varepsilon>0$, there is a unital completely positive map $\theta\colon E\to A$ satisfying that $\|\theta(x)\|\geq(1-\varepsilon)\|x\|$ for $x\in E$, $\|\theta(xy)-\theta(x)\theta(y)\|<\varepsilon$ for those $x,y\in E$ such that $xy\in E$, and $\|\theta(a)-a\|<\varepsilon$ for $a\in E\cap A$. In the $\mathrm{C}^{*}$-probability space setting $(A,\varphi)\subset(B,\psi)$, it is moreover required that $\|\varphi\circ\theta-\psi|_{E}\|<\varepsilon$. It is not hard to see that the class of completely existential embeddings are closed under compositions and tensor products. It is also closed under reduced free products (cf. Corollary 1.9 in [Ro]).
Lemma 10.
For every completely existential embedding $(A,\varphi)\subset(B,\psi)$ of $\mathrm{C}^{*}$-probability spaces and every $(C,\rho)$, the embedding $(A,\varphi)\ast(C,\rho)\subset(B,\psi)\ast(C,\rho)$ is completely existential.
Proof.
Wthout loss of generality, we assume that $B$ is separable. Take an increasing sequence of finite dimensional operator systems $E_{n}\subset B$ such that $\bigcup E_{n}$ is dense in $B$ and $\bigcup(E_{n}\cap A)$ is dense in $A$. Since the embedding $(A,\varphi)\subset(B,\psi)$ is completely existential, for each $n$, there is a unital completely positive map $\theta_{n}\colon E_{n}\to A$ that is approximately isometric, approximately multiplicative, approximately state preserving, and $\theta_{n}|_{E_{n}\cap A}\approx\mathrm{id}_{E_{n}\cap A}$. It extends to a unital completely positive map, still denoted by $\theta_{n}$, from $B$ into $\mathbb{B}(L^{2}(A,\varphi))$. We denote by $\omega$ the vector state that corresponds to the GNS-vector $\hat{1}$ in $L^{2}(A,\varphi)$. One has $\psi_{n}\coloneq\omega\circ\theta_{n}\to\psi$ pointwise on $B$. Let $B_{n}$ denote the quotient of $B$ by the GNS-kernel of $\psi_{n}$. By the Choda–Blanchard–Dykema theorem ([BD] or Theorem 4.8.5 in [BO]), $\theta_{n}\ast\mathrm{id}_{C}$ defines a unital completely positive map from $(B_{n},\psi_{n})\ast(C,\rho)$ into $(\mathbb{B}(L^{2}(A,\varphi)),\omega)\ast(C,\rho)$. By the Skoufranis–Pisier theorem (Theorem 7.1 in [Pi]) for amalgamated free products, the (a priori formally defined) embedding
$(B,\psi)\ast(C,\rho)\hookrightarrow\prod_{n\in\mathbb{N}}(B_{n},\psi_{n})\ast(C,\rho)/\mathcal{U}$
is continuous and locally liftable. By composing this with $\theta_{n}\ast\mathrm{id}_{C}]_{n}$, one obtains a locally liftable embedding of $(B,\psi)\ast(C,\rho)$ into $((\mathbb{B}(L^{2}(A,\varphi)),\omega)\ast(C,\rho))^{\mathcal{U}}$, which lands in $((A,\varphi)\ast(C,\rho))^{\mathcal{U}}$. This witnesses existentialness. ∎
We say a $\mathrm{C}^{*}$-probability space completely selfless if the “complete” analogue of the equivalent conditions in Theorem 2.6 in [Ro] hold. A group is completely $\mathrm{C}^{*}$-selfless if its reduced group $\mathrm{C}^{*}$-algebra is completely selfless. Note that no nuclear tracial $\mathrm{C}^{*}$-probability space is completely selfless, since if it were, then the tracial state on $\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})$ would be amenable (see Chapter 6 in [BO]), which is absurd. Theorems 1 and 2 are upgraded as follows.
Theorem 11.
An infinite countable discrete group having a topologically-free extreme boundary is completely $\mathrm{C}^{*}$-selfless. The tensor product of separable and completely selfless $\mathrm{C}^{*}$-probability spaces is completely selfless. In particular, the class of countable discrete groups that are completely $\mathrm{C}^{*}$-selfless is closed under direct product.
The proof is same as Theorems 1 and 2, but uses the following fact ([Sh] or Example 4.6.11 and Exercise 4.8.1 in [BO]).
Lemma 12.
Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space, $(\mathcal{T},\omega)$ be the Toeplitz $\mathrm{C}^{*}$-probability space, and $C$ be a unital $\mathrm{C}^{*}$-algebra. Then $C\otimes((A,\varphi)\ast(\mathcal{T},\omega))$ is the universal $\mathrm{C}^{*}$-algebra generated by $C\otimes A$ and an isometry $T$ (the generator of the Toeplitz algebra) that satisfies $T^{*}(c\otimes a)T=\varphi(a)(c\otimes 1)$ for $a\in A$ and $c\in C$. Moreover, the free product state $\psi$ on $(A,\varphi)\ast(\mathcal{T},\omega)$ is the unique state that satisfies $\psi|_{A}=\varphi$ and $\psi(aTT^{*}a^{*})=0$ for $a\in A$.
We do not use it, but note the fact that $(A,\varphi)\ast(\mathcal{T},\omega)$ is simple and purely infinite, provided that $A\cap\mathbb{K}(L^{2}(A,\varphi))=0$ ([Ku]).
Theorem 13.
Let $(A,\varphi)$ be a $\mathrm{C}^{*}$-probability space. Suppose that there are a $\mathrm{C}^{*}$-probability space $(B,\psi)$ containing $(A,\varphi)$, an ultrafilter $\mathcal{U}$, and an isometry $T\in B^{\mathcal{U}}$ that satisfies that $T^{*}aT=\varphi(a)$ and $\psi^{\mathcal{U}}(aTT^{*}a^{*})=0$ for all $a\in A$ and that $(T+T^{*})/2\in A^{\mathcal{U}}$. Then $(A,\varphi)$ is completely selfless.
Proof.
Let $(A_{1},\varphi_{1})\subset(A_{2},\varphi_{2})$ denote $(A,\varphi)*(\mathcal{C}_{1},\tau)\subset(A,\varphi)*(\mathcal{T},\omega)$, see Section 5. Let a $\mathrm{C}^{*}$-probability space $(C,\theta)$ be given. We have to show there is an embedding
$(C\otimes A_{1},\theta\otimes\varphi_{1})\hookrightarrow((C\otimes A)^{\mathcal{U}},(\theta\otimes\varphi)^{\mathcal{U}})$
that extends the diagonal embedding of $C\otimes A$. We view $T$ as $1_{C}\otimes T$ in $(C\otimes B)^{\mathcal{U}}$. Then by Lemma 12, the isometry $T$ gives rise to a representation of $C\otimes A_{2}$ in $(C\otimes B)^{\mathcal{U}}$. Since the state $(\theta\otimes\psi)^{\mathcal{U}}$ on $(C\otimes B)^{\mathcal{U}}$ restricts to $\theta\otimes\varphi_{2}$ on $C\otimes A_{2}$, this representation is a state-preserving embedding of $C\otimes A_{2}$ into $(C\otimes B)^{\mathcal{U}}$. Since the generator $(T+T^{*})/2$ of $\mathcal{C}_{1}$ belongs to $A^{\mathcal{U}}$, it restricts to a desired embedding of $C\otimes A_{1}$ into $(C\otimes A)^{\mathcal{U}}$. ∎
7. Proof of Theorem 3
Proof of Theorem 3.
We deal with the first case. By Glimm’s lemma, $\varphi$ is approximated on any finite subset $F$ of $A$ by a pure state $\psi$ that are disjoint from $\varphi$. By excision (see Proposition 11.4.2 in [BO]), there is a net of projections $p_{j}$ such that $\|p_{j}ap_{j}-\psi(a)p_{j}\|\to 0$ for $a\in A$. The net $(p_{j})_{j}$ converges in $A^{**}$ to the support projection of the pure state $\psi$. Hence $\varphi(ap_{j}a^{*})\to 0$ for $a\in A$. Thus, by pure infiniteness, there are a directed set $I$ and a net $(T_{i})_{i\in I}$ of isometries in $A$ such that $\|T_{i}^{*}aT_{i}-\varphi(a)\|\to 0$ and $\varphi(aT_{i}T_{i}^{*}a^{*})\to 0$ for $a\in A$. Let $\mathcal{U}$ be a cofinal ultrafilter on $I$. The isometry $T\coloneq[T_{i}]_{i}$ in $A^{\mathcal{U}}$ satisfies $T^{*}aT=\varphi(a)$ and $\varphi^{\mathcal{U}}(aTT^{*}a^{*})=0$ for $a\in A$. By Theorem 13, we are done.
We deal with the second case. Since $A$ is $\mathcal{Z}$-stable and $\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})\hookrightarrow\mathcal{Z}^{\mathcal{U}}$ (Theorem 4.1 in [Oz]), the embedding $A\subset A\otimes C^{*}_{\lambda}(\mathbf{F}_{\infty})$ is existential by exactness, where we identify $A$ with $A\otimes\mathbb{C}1$. Since $A$ is simple and monotracial, the uniform Dixmier property holds by the Haagerup–Zsido theorem ([HZ]). Namely, there is a net $(F_{i})_{i\in I}$ of finite families of unitary elements in $A$ that satisfies $|F_{i}|^{-1}\sum_{u\in F_{i}}uau^{*}\to\tau(a)$ for $a\in A$. By replacing $F_{i}$ with $F_{i}^{*}F_{i}$, we may assume that the finite set $F_{i}$ is closed under the adjoint operation. Let $\{u^{(i)}_{n}:n=1,\ldots,|F_{i}|\}$ be an enumeration of $F_{i}$. Since $A\subset A\otimes\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})$ is existential, by replacing $u^{(i)}_{n}$ with likenesses in $A$ of $u^{(i)}_{n}\otimes t_{n}$ in $A\otimes\mathrm{C}^{*}_{\lambda}(\mathbf{F}_{\infty})$, we may further assume by Voiculescu’s inequality (see Proof of Theorem 2) that $\||F_{i}|^{-1}\sum_{u\in F_{i}}uau\|\to 0$ for $a\in A$. Let $\mathcal{U}$ be a cofinal ultrafilter on $I$.
Let’s consider the Cuntz algebra $(\mathcal{O}_{\infty},\omega)$ generated by isometries $\{l_{n}\}$ with mutually orthogonal ranges, together with the vacuum state. Recall that $\{(l_{n}+l_{n}^{*})/2\}$ generates the free semicircular system $(\mathcal{C},\tau)\subset(\mathcal{O}_{\infty},\omega)$ (see [VDN]). The elements
$T_{i}\coloneq(2|F_{i}|)^{-1/2}\sum_{n}(u^{(i)}_{n}+(u^{(i)}_{n})^{*})\otimes l_{n}$
in $A\otimes\mathcal{O}_{\infty}$ satisfies $T_{i}^{*}aT_{i}\to\tau(a)$ and $(\tau\otimes\omega)(aT_{i}T_{i}^{*}a^{*})=0$ for $a\in A$. Moreover, $(T_{i}+T_{i}^{*})/2\in A\otimes\mathcal{C}$. Hence by Lemma 12, it gives rise to an embedding of $(A,\tau)\ast(\mathcal{T},\omega)$ into $(A\otimes\mathcal{O}_{\infty},\tau\otimes\omega)^{\mathcal{U}}$. It restricts to an embedding of $(A,\tau)\ast(\mathcal{C}_{1},\tau)$ into $(A\otimes\mathcal{C},\tau\otimes\tau)^{\mathcal{U}}$ that witnesses the relative existentialness (see [Ro]) of $(A,\tau)\hookrightarrow(A,\tau)\ast(\mathcal{C}_{1},\tau)$ in $(A\otimes\mathcal{C},\tau\otimes\tau)$. Since the embedding of $(A,\tau)$ into $(A\otimes\mathcal{C},\tau\otimes\tau)$ is existential by assumption, it follows that the embedding of $(A,\tau)$ into $(A,\tau)\ast(\mathcal{C}_{1},\tau)$ is also existential. ∎
8. The Powers property and selflessness
$\mathrm{C}^{*}$-simplicity of the free group of rank $2$ has been proved by R. T. Powers ([Po]) by a combinatorial method. The combinatorial method is later formalized by P. de la Harpe ([dlH]) as the Powers property. The following is very close to it in spirit. It is designed to allow the Haagerup–Pisier type decomposition ([HP]).
Definition.
We say that a group $\Gamma$ has property $\mathrm{P}_{\mathrm{PHP}}$ if for every finite subset $F\subset\Gamma$ and $\varepsilon>0$, there are $n\in\mathbb{N}$ and $t_{i}\in\Gamma$ and $C_{i}\subset D_{i}\subset\Gamma$ for $1\leq i\leq n$ satisfying that the members of
$\{aC_{i}:a\in F,\,i=1,\ldots,n\}\cup\{bt_{j}^{-1}(\Gamma\setminus D_{j}):b\in F,\,j=1,\ldots,n\}$
are mutually disjoint and that
$\sup_{x\in\Gamma}|\{i:x\in D_{i}\}\cup\{j:x\in t_{j}^{-1}(\Gamma\setminus C_{j})\}|\leq\varepsilon n^{1/2}.$
It is not too hard to see that the class of groups with $\mathrm{P}_{\mathrm{PHP}}$ is closed under the direct product; for $t_{i}^{(k)}\in\Gamma^{(k)}$ and $C_{i}^{(k)}\subset D_{i}^{(k)}\subset\Gamma^{(k)}$, $k=1,2$, set $t_{i}\coloneq(t_{i}^{(1)},t_{i}^{(2)})\in\Gamma^{(1)}\times\Gamma^{(2)}$, $C_{i}\coloneq C_{i}^{(1)}\times C_{i}^{(2)}$, and $D_{i}\coloneq(D_{i}^{(1)}\times\Gamma^{(2)})\cup(\Gamma^{(1)}\times D_{i}^{(2)})$.
Theorem 14.
A group $\Gamma$ with property $\mathrm{P}_{\mathrm{PHP}}$ is completely $\mathrm{C}^{*}$-selfless.
Proof.
Given a finite subset $F\subset\Gamma$ and $\varepsilon>0$, we take $t_{i}$ and $C_{i}\subset D_{i}\subset\Gamma$ for $1\leq i\leq n$ as in Definition. Let $P_{i}\leq Q_{i}$ denote the diagonal projections in $\mathbb{B}(\ell_{2}\Gamma)$ corresponding to the subsets $C_{i}\subset D_{i}$ and set $P_{i}^{\prime}\coloneq t_{i}^{-1}Q_{i}^{\perp}t_{i}$ and $Q_{i}^{\prime}\coloneq t_{i}^{-1}P_{i}^{\perp}t_{i}$. Note that $\{P_{i}\}\cup\{P^{\prime}_{j}\}$ are mutually orthogonal. Here $\Gamma$ is represented in $\mathbb{B}(\ell_{2}\Gamma)$ by the left regular representation. We denote by $\omega$ the vector state on $\mathbb{B}(\ell_{2}\Gamma)$ corresponding to $\delta_{1}$, which satisfies $\omega|_{\mathrm{C}^{*}_{\lambda}(\Gamma)}=\tau$. We define $T\in\mathbb{B}(\ell_{2}\Gamma)$ by
$T\coloneq(2n)^{-1/2}\sum_{i}(P_{i}t_{i}+t_{i}^{-1}Q_{i}^{\perp})=(2n)^{-1/2}\sum_{i}(P_{i}t_{i}+P^{\prime}_{i}t_{i}^{-1}).$
The element $T$ is an almost isometry as
$T^{*}T=(2n)^{-1}\sum_{i}(t_{i}^{-1}P_{i}t_{i}+t_{i}P_{i}^{\prime}t_{i}^{-1})=1-(2n)^{-1}\sum_{i}(Q_{i}^{\prime}+Q_{i})\approx_{\varepsilon n^{-1/2}}1.$
Moreover, $T^{*}aT=0=\tau(a)$ for every $a\in F\setminus\{1\}$. One has by Schur’s test
$\|\sum_{i}(Q_{i}-P_{i})t_{i}\|$ $=\|\sum_{i}(Q_{i}-P_{i})t_{i}(Q_{i}^{\prime}-P_{i}^{\prime})\|\leq\|\sum_{i}Q_{i}\|^{1/2}\|\sum_{i}Q^{\prime}_{i}\|^{1/2}\leq\varepsilon n^{1/2}$
and hence
$\|T+T^{*}-(2n)^{-1/2}\sum_{i}(t_{i}+t_{i}^{*})\|\leq 2(2n)^{-1/2}\|\sum_{i}(Q_{i}-P_{i})t_{i}\|\leq 2\varepsilon.$
Also, for every finite subset of $\Gamma$, the corresponding diagonal projection $R\in\mathbb{B}(\ell_{2}\Gamma)$ satisfies $\|RT\|^{2}\leq(2n)^{-1}\|\sum_{i}P_{i}+P_{i}^{\prime}\|\|\sum_{i}t_{i}^{-1}Rt_{i}+t_{i}Rt_{i}^{-1}\|\leq n^{-1}\operatorname{rank}R$. It follows that a suitable ultra-limit $\tilde{T}\in\mathbb{B}(\ell_{2}\Gamma)^{\mathcal{U}}$ of $T$ as above verifies the condition of Theorem 13 for $(\mathrm{C}^{*}_{\lambda}(\Gamma),\tau)\subset(\mathbb{B}(\ell_{2}\Gamma),\omega)$ and hence $(\mathrm{C}^{*}_{\lambda}(\Gamma),\tau)$ is completely selfless. ∎
Proposition 15.
Let $\Gamma$ be a group. Assume that there is a continuous action of $\Gamma$ on a Hausdorff topological space that is minimal, extremely proximal, and topologically free. Then $\Gamma$ has property $\mathrm{P}_{\mathrm{PHP}}$.
Proof.
We sketch the proof by adopting that of Lemma 4 in [dlH]. Let a finite subset $F\subset\Gamma$ and $\varepsilon>0$ be given. Take $n\geq\varepsilon^{-2}$. Then, there are $t_{1},\ldots,t_{n}$ in $\Gamma$ and mutually disjoint open subsets $U_{i}^{\pm}$ such that $t_{i}(L\setminus U_{i}^{-})\subset U_{i}^{+}$ for every $i$. Moreover, we may assume that $a(U_{i}^{+}\cup U_{i}^{-})\cap(U_{j}^{+}\cup U_{j}^{-})=\emptyset$ for every $a\in F\setminus\{1\}$ and every $i,j$. We fix a point $x_{0}\in L$ and set $C_{i}\coloneq D_{i}\coloneq\{s\in\Gamma:sx_{0}\in U_{i}^{+}\}$. They witness $\mathrm{P}_{\mathrm{PHP}}$. ∎
Acylindrically hyperbolic groups with trivial finite radical satisfy the assumption of Proposition 15 (Proposition 0.3 in [AD]). Hence it removes property RD assumption from Amrutam et al.’s theorem ([AGKEP]), see also [Ya] for an alternative proof. However, unlike Amrutam et al. and Yang’s theorem ([AGKEP] [Ya]), it does not provide an element in $\Gamma^{\mathcal{U}}$ that is strongly free from $\Gamma$. The same comment applies to Theorem 1 and the following proposition that partly generalizes Vigdorovich’s theorem ([Vi]).
Proposition 16.
Every Zariski-dense subgroup $\Gamma\leq\mathrm{PSL}(d\geq 2,\mathbb{R})$ satisfies property $\mathrm{P}_{\mathrm{PHP}}$.
Proof.
We follow the proof of Proposition 1 in [BCH]. Let $B:=\mathbb{P}^{d-1}(\mathbb{R})$ denote the homogeneous space of $G\coloneq\mathrm{PSL}(d,\mathbb{R})$, on which $\Gamma$ acts continuously. Take $t_{0}\in\Gamma$ with eigenvalues $\lambda_{1}>\lambda_{2}>\cdots>\lambda_{n}>0$ and denote by $\xi_{1}:\cdots:\xi_{n}]$ the corresponding homogeneous coordinates on $B$. We set $p\coloneq[1:0:\cdots:0]\in K\coloneq\{[\xi_{1}:\cdots:\xi_{n}]:\xi_{n}=0\}$ and $p^{\prime}\coloneq[0:\cdots:0:1]\in K^{\prime}\coloneq\{[\xi_{1}:\cdots:\xi_{n}]:\xi_{1}=0\}$. Then for every $x\in B\setminus\{p\}$, every limit point of $t_{0}^{-n}x$ is in $K^{\prime}$; and for every $x\in B\setminus K$, one has $\lim_{n}t_{0}^{-n}x=p^{\prime}$.
Let a finite subset $F\subset\Gamma$ and $\varepsilon>0$ be given. We claim that for every $n$ there are $s_{1},\ldots,s_{n}\in\Gamma$ such that the members of
$\{as_{i}p:a\in F,i=1,\ldots,n\}\cup\{bs_{j}p^{\prime}:b\in F,j=1,\ldots,n\}$
are mutually distinct and that the hyperplanes $\{s_{i}K\}$ (resp. $\{s_{i}K^{\prime}\}$) are in general position, i.e., for every $I\subset\{1,\ldots,n\}$ the linear dimension of $\bigcap_{i\in I}s_{i}K$ is $\max\{d-|I|,0\}$ (resp. likewise). The proof is by induction on $n$. Suppose that the claim holds true for n1n-1 and we are given $s_{1},\ldots,s_{n-1}\in\Gamma$ that meet the claim. For each $a\in\Gamma$ and $q,q^{\prime}\in B$ (which may be the same), the subset $\{s\in\Gamma:sq\neq aq^{\prime}\}$ is non-empty and Zariski-open. For each $a\in\Gamma\setminus\{1\}$ and $q,q^{\prime}\in B$, the subset $\{s\in\Gamma:sq\neq asq^{\prime}\}$ is also non-empty and Zariski-open. For example, suppose by contradiction that $sq=asq^{\prime}$ for all $s\in\Gamma$. Then $q=aq^{\prime}$ and $sq=asq^{\prime}=asa^{-1}q$ for all $s\in G$ by Zariski-density of $\Gamma$ in $G$, but since $\bigcup_{s}s^{-1}G_{q}s=G$, where $G_{q}$ is the stabilizers of $q$, this implies that $a$ acts as identity on $B$, meaning that $a=1$. Also for each proper subspaces $L$ and $L^{\prime}$ the subset $\{s\in\Gamma:L^{\prime}\not\subset sL\}$ is non-empty and Zariski-open. Since a finite intersection of non-empty Zariski-open subsets is non-empty, there is $s_{n}$ such that $s_{1},\ldots,s_{n}$ meet the claim.
We continue the proof and let $s_{1},\ldots,s_{n}$ be as in the previous claim. We choose neighborhoods $U_{i}$ of $s_{i}p$ and $V_{i}$ of $s_{i}K$ with $U_{i}\subset V_{i}$ and likewise $U_{i}^{\prime}\subset V_{i}^{\prime}$ that satisfy that
$\{aU_{i}:a\in F,i=1,\ldots,n\}\cup\{bU_{j}^{\prime}:b\in F,j=1,\ldots,n\}$
are mutually disjoint and that every $I\subset\{1,\ldots,n\}$ with cardinality $d$ satisfies $\bigcap_{i\in I}V_{i}=\emptyset$ and $\bigcap_{i\in I}V_{i}^{\prime}=\emptyset$. Then by compactness, for each $i$ there is $n_{i}\in\mathbb{N}$ large enough so that $t_{i}\coloneq s_{i}t_{0}^{n_{i}}s_{i}^{-1}$ satisfies $t_{i}^{-1}(B\setminus U_{i})\subset V_{i}^{\prime}$ and $t_{i}^{-1}(B\setminus V_{i})\subset U_{i}^{\prime}$. We further choose $x_{0}\in B$ and set $C_{i}\coloneq\{g\in\Gamma:gx_{0}\in U_{i}\}$ and $D_{i}\coloneq\{g\in\Gamma:gx_{0}\in V_{i}\}$. Thus, for $n$ large enough, $t_{1},\ldots,t_{n}$ and $C_{i}\subset D_{i}$ verify $\mathrm{P}_{\mathrm{PHP}}$ with
$\sup_{x\in\Gamma}|\{i:x\in D_{i}\}\cup\{j:x\in t_{j}^{-1}(\Gamma\setminus C_{j})\}|\leq 2(d-1)\leq\varepsilon n^{1/2}.$